# When Parts Are Greater Than Sums: Individual LLM Components Can Outperform Full Models

### v2

@@ 5.1 Proposed Method
Our findings in section 4 show the promising direction of selecting internal components to improve ICL. Therefore, we propose a method that reweighs components by learning a weight $w_{j}\in\mathbb{R}$ on every component activation $C_{j}$. Reweighting is a soft version of selection, which can be learned by gradient descent on very few examples.
@@
where $U_{\mathcal{Y}}\in\mathbb{R}^{|\mathcal{Y}|\times d}$ is a submatrix of $U$ that comprises the output embeddings of label words, $P_{rw}$ is the probability distribution of the LLM after reweighting, and $\lambda$ is the hyperparameter of the $L_{1}$ loss to encourage sparsity on the component weights. We obtain the activations $\{C_{j}\}_{j=1}^{N}$ of all components in one $K^{\prime}$-shot forward pass, computed on the prompt derived from $\mathcal{D}_{\text{demo}}$, followed by $x$. Our method scales each component’s direct contributions to the logits ($U_{\mathcal{Y}}\cdot C_{j}\in\mathbb{R}^{|\mathcal{Y}|}$) by $w_{j}$. In practice, we cache these contributions on the entire training set as input features to the linear layer $w$, which allows us to discard the entire LLM while training $w$ (line 9 and 13 in Algorithm 1), saving tremendous training time and GPU memory. The cache only requires $O(|\mathcal{Y}|\times N\times|\mathcal{D}_{\text{train}}|)$ space. At inference, the overhead of our method over $K^{\prime}$-shot ICL is to early decode $N$ components and apply the learned weights, i.e., $\sum_{j=1}^{N}w_{j}\;(U_{\mathcal{Y}}\cdot C_{j})$. As both $|\mathcal{Y}|$ and $N$ are small333$|\mathcal{Y}|\leq 4,N<2000$ for all tasks and LLMs in this paper., the overhead is negligible compared to the computation of the LLM itself.
Algorithm 1 Component Reweighting
5:Initialize $\mathcal{G}^{\text{dev}}\leftarrow\varnothing$
6:for  $(x,y)\in\mathcal{D}_{\text{train}}$  do
8:     for  $j\leftarrow 1$ to $N$  do
9:         $\mathcal{G}^{\text{dev}}\leftarrow\mathcal{G}^{\text{dev}}\cup(U_{\mathcal{Y}}\cdot C_{j})$triangleright early decode
10:     end for
11:end for
12:Initialize $w\leftarrow[1,\dotsc,1]\in\mathbb{R}^{N}$
13:Train the weights $w$ on $\mathcal{G}^{\text{dev}}$ with Eq. 9
15:for  $(x,y)\in\mathcal{D}_{\text{test}}$  do
18:     for  $j\leftarrow 1$ to $N$  do $\triangleright$ Test-Time Overhead
20:     end for
21:     $\hat{y}\leftarrow{\arg\max}_{y\in\mathcal{Y}}\;\mathbf{g}$
22:     $\mathcal{Z}\ \leftarrow\mathcal{Z}\ \cup\hat{y}$
23:end for
24:return $\mathcal{Z}$
5.2 Baselines
Standard ICL.
Prompt Selection.
Calibration.
As LLMs tend to predict a certain class over others, Zhao et al. (2021) reweight the output class probabilities. They use context-free inputs, such as “N/A”, to calibrate the probability distribution. However, Fei et al. (2023); Zhou et al. (2023) find context-free inputs sometimes ineffective, because in-domain context is important for calibration. Thus, we introduce Calib+, which calibrates the original probabilities $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ with a training set of in-distribution labeled examples, $\mathcal{D}_{\text{train}}$. We train the calibration weights $\mathbf{v}\in\mathbb{R}^{|\mathcal{Y}|}$ on $\mathcal{D}_{\text{train}}$ with cross-entropy loss and obtain the calibrated probabilities $\hat{\mathbf{p}}=\text{softmax}(\mathbf{v}\cdot\mathbf{p})$. For direct comparisons, we split the $K$ examples into the same $\mathcal{D}_{\text{demo}}$ and $\mathcal{D}_{\text{train}}$ sets as component reweighting for Calib+, where $|\mathcal{D}_{\text{demo}}|=K^{\prime}$. We include the training details of both methods in A.7.
@@ 5.3 Results
We set $K=\{12,24\}$ following common practices. Table 4 compares our component reweighting (CompRW) with standard ICL (Standard), prompt selection (PromptS), and calibration (Calib+). First, we find that simply increasing the number of demonstrations from $4$ to $24$ has limited improvements in ICL accuracy, while the longer prompt greatly increases the inference time. For example, on Llama-2-7B, Standard $24$ only improves the average accuracy by $2.6\%$ over Standard $3,4$ and the accuracy even decreases on Mistral-Instruct. Second, PromptS performs the worst in most setups, likely because it is hard to find similar examples from a small pool of $K$ examples, and a bad selection induces majority label biases. Third, both calibration (Calib+) and component reweighting (CompRW) achieve substantially better accuracy than Standard $3,4$ with little test-time overhead. Overall, CompRW achieves the best average accuracy in all setups, outperforming Standard $12$ by $6.0\%,1.8\%,2.6\%$ on Llama-2-7B, Llama-2-13B, and Mistral-Instruct-7B, respectively, and outperforming Standard $24$ by $6.0\%,2.2\%,5.1\%$. We run one-tailed paired t-tests comparing CompRW with Calib+ and find that p-values $<0.05$ in all 6 setups (see Table 5), showing that CompRW performs significantly better Calib+.
@@ 6 When Do Good Components Emerge?
We study the dynamics of components during pretraining by monitoring their accuracies on 32 checkpoints of Pythia-6.9B, uniformly sampled from the first to the last checkpoint. For each checkpoint, we run 4-shot ICL on AGNews with 3 templates $\times$ 3 sets of demonstrations. Figure 3 shows the average accuracy shaded by the standard deviation.
While the full model (green) fluctuates and has a large variance across prompts, the top-1 components (solid blue) achieve good accuracy at an early step and plateaus quickly. We also backtrack the top-1 components of different prompts at the last checkpoint (dashed blue), monitoring how they perform on average during pretrain. We observe that they are not the top components at the early stage (there are gaps between the two blue lines before the $75k$ steps), but start to perform steadily well from the middle stage. Our findings also hold on SST2 and Pythia-1.4B (Figure 6 in the appendix), suggesting that the model’s ability to do a task emerges before it is apparent from the full model.
7 Related Work
@@
Prior work shows that ICL performance varies greatly across different choices of demonstrations and templates Zhao et al. (2021); Lu et al. (2022); Sclar et al. (2024); Voronov et al. (2024). While several approaches, such as prompt selection Liu et al. (2022b); Chang and Jia (2023); Fu et al. (2023), prompt ensemble Min et al. (2022a); Arora et al. (2023); Voronov et al. (2024), and many-shot ICL Agarwal et al. (2024), address this issue, they treat LLMs like black boxes without studying the internals. Besides, they greatly increase inference time or require a large set of labeled data, which deviates from true few-shot learning Perez et al. (2021). In comparison, our paper studies this problem by looking inside the LLMs. Rather than selecting prompts, we select components in a soft, learnable way. Our method only requires $\{12,24\}$ examples and has little computation overhead over $4$-shot ICL at inference.
@@
Components interpretation studies the function of different components in a trained model Csordás et al. (2021); Elhage et al. (2021); Shah et al. (2024), where components could be neurons Radford et al. (2017); Wang et al. (2022a); Gurnee et al. (2023), attention heads Olsson et al. (2022), and MLPs Geva et al. (2021). To analyze the components, probing Alain and Bengio (2017), ablation Wang et al. (2023a, b); Chang et al. (2024), patching Wang et al. (2023a); Goldowsky-Dill et al. (2023), and early decoding nostalgebraist (2020); Geva et al. (2022) are widely used techniques. For example, Li et al. (2024) train a linear probe on every attention head to discover the truthful heads inside LLMs. Michel et al. (2019); Voita et al. (2019) prune away a large percentage of attention heads and show that only a few are critical to the performance. Hendel et al. (2023); Liu et al. (2023); Merullo et al. (2024b); Todd et al. (2024) view ICL as compressing demonstrations into function vectors, where they remove the demonstrations and modify (patch) the LLM activations at certain layers with the function vectors at test time. Early decoding interprets the investigated components in the textual space by projecting them through the output embedding matrix Geva et al. (2023). Our model decomposition is based on early decoding and we share some similarities with prior work Yu et al. (2023); Wang et al. (2023c); Merullo et al. (2024a), especially in discovering individual components that perform well on a task. Our contributions lie in providing a new view of ICL by decomposition, which reveals the transferability of components across diverse ICL settings.
@@ 9 Limitations
Our component reweighting method requires a small set of labeled data $\mathcal{D}_{\text{train}}$ to train the component weights $w$. However, we believe it is not unreasonable to have at least $K=12$ labeled examples in total and we compare with baselines using the same $K$ examples. On the other hand, we do not compare with fine-tuning-based baselines, such as LM-BFFGao et al. (2021), T-few Liu et al. (2022a), and LoRA Hu et al. (2021), because they usually require a larger GPU memory for training and more sophisticated early stopping criteria to prevent overfitting on few-shot examples. Another limitation is that we only experiment with classification tasks for easy evaluation. We leave it for future work to generalize our method to generation tasks by doing decomposition and reweighting at every token during generation.
@@ Acknowledgements
We thank Johnny Wei, Qinyuan Ye, Ryan Wang, Gustavo Lucas Carvalho, Ameya Godbole, Wang Zhu, and Daniel Firebanks-Quevedo for their helpful feedback. This work was funded in part by gifts from Open Philanthropy and Cisco Research.
@@ A.1 Tasks and Templates
Table 8 summarizes the 13 datasets we use in the paper, where we construct balanced test sets by randomly sampling 2000 examples in each task. We form the prompts by concatenating demonstrations in a randomly shuffled order. To avoid the recency bias Zhao et al. (2021), we keep shuffling the demonstrations until the last two have different labels. For minimally conservative templates (section 4.1), Table 9 compares the contrast sets we construct on Llama-2-7B. For our case study on Task069 and Task070, we sample 3 templates from Sclar et al. (2024). In section 4.3, we calculate the correlation between component accuracies of every pair of runs from the two tasks under a fixed template. The correlation score ($r=-0.6$) is averaged across runs and templates. Figure 8 compare the prompts of Task069 and Task070, which consist of an instruction followed by $K$ templated demonstrations. Originally, the two tasks have $\sim 4\%$ of parallel examples. To make our task transfer challenging, we discard these overlapped examples.

### v3

@@ 5.1 Proposed Method
Our findings in section 4 show the promising direction of selecting internal components to improve ICL. Therefore, we propose a method that reweights components by learning a weight $w_{j}\in\mathbb{R}$ on every component activation $C_{j}$. Reweighting is a soft version of selection, which can be learned by gradient descent on very few examples.
where $U_{\mathcal{Y}}\in\mathbb{R}^{|\mathcal{Y}|\times d}$ is a submatrix of $U$ that comprises the output embeddings of label words, $P_{rw}$ is the probability distribution of the LLM after reweighting, and $\lambda$ is the hyperparameter of the $L_{1}$ loss to encourage sparsity on the component weights. We obtain the activations $\{C_{j}\}_{j=1}^{N}$ of all components in one $K^{\prime}$-shot forward pass, computed on the prompt derived from $\mathcal{D}_{\text{demo}}$, followed by $x$. Our method scales each component’s direct contributions to the logits ($U_{\mathcal{Y}}\cdot C_{j}\in\mathbb{R}^{|\mathcal{Y}|}$) by $w_{j}$. In practice, we cache these contributions on the entire training set as input features to the linear layer $w$, which allows us to discard the entire LLM while training $w$ (line 9 and 13 in Algorithm 1), saving tremendous training time and GPU memory. The cache only requires $O(|\mathcal{Y}|\times N\times|\mathcal{D}_{\text{train}}|)$ space. At inference time, the overhead of our method over $K^{\prime}$-shot ICL is to early decode $N$ components and apply the learned weights, i.e., $\sum_{j=1}^{N}w_{j}\;(U_{\mathcal{Y}}\cdot C_{j})$. As both $|\mathcal{Y}|$ and $N$ are small ($N<2000$ for all LLMs in this paper), the overhead is negligible compared to the computation of the LLM itself.
Algorithm 1 Component Reweighting
5:Initialize $\mathcal{G}^{\text{train}}\leftarrow\varnothing$
6:for  $(x,y)\in\mathcal{D}_{\text{train}}$  do
8:     for  $j\leftarrow 1$ to $N$  do
9:         $\mathcal{G}^{\text{train}}\leftarrow\mathcal{G}^{\text{train}}\cup(U_{\mathcal{Y}}\cdot C_{j})$triangleright early decode
10:     end for
11:end for
12:Initialize $w\leftarrow[1,\dotsc,1]\in\mathbb{R}^{N}$
13:Train the weights $w$ on $\mathcal{G}^{\text{train}}$ with Eq. 9
15:for  $(x,y)\in\mathcal{D}_{\text{test}}$  do
18:     for  $j\leftarrow 1$ to $N$  do $\triangleright$ Test-Time Overhead
20:     end for
21:     $\hat{y}\leftarrow{\arg\max}_{y\in\mathcal{Y}}\;\mathbf{g}$
22:     $\mathcal{Z}\ \leftarrow\mathcal{Z}\ \cup\hat{y}$
23:end for
24:return $\mathcal{Z}$
5.2 Baselines
Standard ICL.
Prompt Selection.
Calibration.
As LLMs tend to predict a certain class over others, Zhao et al. (2021) reweight the output class probabilities. They use context-free inputs, such as “N/A”, to calibrate the probability distribution. However, Fei et al. (2023) and Zhou et al. (2023) find context-free inputs sometimes ineffective, because in-domain context is important for calibration. Thus, we introduce Calib+, which calibrates the original probabilities $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ with a training set of in-distribution labeled examples, $\mathcal{D}_{\text{train}}$. We train the calibration weights $\mathbf{v}\in\mathbb{R}^{|\mathcal{Y}|}$ on $\mathcal{D}_{\text{train}}$ with cross-entropy loss and obtain the calibrated probabilities $\hat{\mathbf{p}}=\text{softmax}(\mathbf{v}\cdot\mathbf{p})$. For direct comparisons, we split the $K$ examples into the same $\mathcal{D}_{\text{demo}}$ and $\mathcal{D}_{\text{train}}$ sets as component reweighting for Calib+, where $|\mathcal{D}_{\text{demo}}|=K^{\prime}$. We include the training details of both methods in A.8.
@@ 5.2 Baselines
Llama-3-8B Standard $3,4$ $91.4_{\hskip 1.42271pt1.7}$ $79.2_{\hskip 1.42271pt7.2}$ $74.0_{\hskip 1.42271pt8.0}$ $58.7_{\hskip 1.42271pt4.7}$ $76.5_{\hskip 1.42271pt2.2}$ $59.4_{\hskip 1.42271pt3.7}$ $84.0_{\hskip 1.42271pt6.6}$ $87.4_{\phantom{1}\hskip 1.42271pt5.5}$ $76.3$
Standard $12$ $\textbf{92.2}_{\hskip 1.42271pt0.6}$ $\textbf{79.6}_{\hskip 1.42271pt6.9}$ $73.1_{\hskip 1.42271pt5.0}$ $\textbf{63.3}_{\hskip 1.42271pt2.4}$ $77.5_{\hskip 1.42271pt2.2}$ $62.7_{\hskip 1.42271pt3.8}$ $87.7_{\hskip 1.42271pt2.0}$ $82.1_{\hskip 1.42271pt18.1}$ $77.3$
Calib12121212 $91.1_{\hskip 1.42271pt1.5}$ $79.2_{\hskip 1.42271pt5.8}$ $\textbf{77.9}_{\hskip 1.42271pt3.3}$ $60.5_{\hskip 1.42271pt9.3}$ $77.3_{\hskip 1.42271pt2.1}$ $65.2_{\hskip 1.42271pt3.2}$ $86.4_{\hskip 1.42271pt4.7}$ $\textbf{87.7}_{\phantom{1}\hskip 1.42271pt4.0}$ $78.2$
CompRW $12$ $90.7_{\hskip 1.42271pt2.0}$ $78.3_{\hskip 1.42271pt6.7}$ $77.2_{\hskip 1.42271pt2.9}$ $61.8_{\hskip 1.42271pt6.4}$ $\textbf{78.0}_{\hskip 1.42271pt1.8}$ $\textbf{66.9}_{\hskip 1.42271pt2.4}$ $\textbf{89.1}_{\hskip 1.42271pt1.0}$ $87.4_{\phantom{1}\hskip 1.42271pt3.8}$ 78.7
Standard $24$ $\textbf{92.2}_{\hskip 1.42271pt0.8}$ $78.2_{\hskip 1.42271pt7.2}$ $78.0_{\hskip 1.42271pt2.0}$ $\textbf{63.8}_{\hskip 1.42271pt1.8}$ $76.2_{\hskip 1.42271pt3.0}$ $66.4_{\hskip 1.42271pt2.3}$ $87.9_{\hskip 1.42271pt1.9}$ $80.6_{\hskip 1.42271pt18.9}$ $77.9$
Calib24242424 $91.7_{\hskip 1.42271pt1.4}$ $\textbf{80.0}_{\hskip 1.42271pt6.1}$ $78.3_{\hskip 1.42271pt3.4}$ $\textbf{63.8}_{\hskip 1.42271pt2.9}$ $78.1_{\hskip 1.42271pt1.7}$ $66.0_{\hskip 1.42271pt2.4}$ $86.7_{\hskip 1.42271pt4.9}$ $\textbf{87.7}_{\phantom{1}\hskip 1.42271pt3.8}$ $79.0$
CompRW $24$ $91.6_{\hskip 1.42271pt1.7}$ $79.1_{\hskip 1.42271pt6.9}$ $\textbf{78.8}_{\hskip 1.42271pt2.7}$ $63.7_{\hskip 1.42271pt3.3}$ $\textbf{78.5}_{\hskip 1.42271pt1.4}$ $\textbf{67.4}_{\hskip 1.42271pt2.6}$ $\textbf{89.5}_{\hskip 1.42271pt1.1}$ $87.4_{\phantom{1}\hskip 1.42271pt3.2}$ 79.5
@@ 5.3 Results
We set $K=\{12,24\}$. Table 4 compares our component reweighting (CompRW) with standard ICL (Standard), prompt selection (PromptS), and calibration (Calib+). First, we find that simply increasing the number of demonstrations from $4$ to $24$ has limited improvements in ICL accuracy, while the longer prompt greatly increases the inference time. For example, on Llama-2-7B, Standard $24$ only improves the average accuracy by $2.6\%$ over Standard $3,4$ and the accuracy even decreases on Mistral-Instruct. Second, PromptS performs the worst in most setups, likely because it is hard to find similar examples from a small pool of $K$ examples, and a bad selection induces majority label biases. Third, both calibration (Calib+) and component reweighting (CompRW) achieve substantially better accuracy than Standard $3,4$ with little test-time overhead. Overall, CompRW achieves the best average accuracy in all setups, outperforming Standard $12$ by $6.0\%,1.8\%,2.6\%,1.4$ on Llama-2-7B, Llama-2-13B, Mistral-Instruct-7B, and Llama-3-8B, respectively, and outperforming Standard $24$ by $6.0\%,2.2\%,5.1\%,1.6\%$, respectively. We run one-tailed paired t-tests comparing CompRW with Calib+ and find that p-values $<0.05$ in all 8 setups (see Table 6), showing that CompRW performs significantly better Calib+.
@@ 6 When Do Good Components Emerge?
We study the dynamics of components during pretraining by monitoring their accuracies on 32 checkpoints of Pythia-6.9B, uniformly spaced from the first to the last checkpoint. For each checkpoint, we run 4-shot ICL on AGNews with 3 templates $\times$ 3 sets of demonstrations. The demonstrations are balanced in labels with randomly shuffled orders. Figure 3 shows the average accuracy of the 9 runs shaded by the standard deviation.
While the full model (green) fluctuates and has a large variance across prompts, the top-1 components (solid blue) achieve good accuracy at an early step and plateau quickly. We also backtrack the top-1 components of different prompts at the last checkpoint (dashed blue), monitoring how they perform on average during pretraining. We observe that they are not the top components at the early stage (there are gaps between the two blue lines before the $75k$ steps), but start to perform steadily well from the middle stage. Our findings also hold on SST2 and Pythia-1.4B (see Figure 6 in the appendix), suggesting that the model’s ability to do a task emerges before it is apparent from the full model on these tasks. [^4] On the other hand, Pythia models perform poorly on the other tasks over all checkpoints; thus, the training dynamics of the model components on challenging tasks remain unclear.
7 Related Work and Discussion
@@
Prior work shows that ICL performance varies greatly across different choices of demonstrations and templates Zhao et al. (2021); Lu et al. (2022). Specifically, Sclar et al. (2024) and Voronov et al. (2024) find no universally better prompt template that can transfer across tasks and models, implying that it is not easy to explain ICL through prompt engineering. While several approaches, such as prompt selection Liu et al. (2022b); Chang and Jia (2023); Fu et al. (2023), prompt ensemble Min et al. (2022a); Arora et al. (2023); Voronov et al. (2024), and many-shot ICL Agarwal et al. (2024), substantially improve accuracy, they treat LLMs like black boxes without understanding the internals. Besides, they greatly increase inference time or require a large set of labeled data, which deviates from true few-shot learning Perez et al. (2021). In comparison, our paper studies this problem by looking inside the LLMs. Rather than selecting prompts, we select components in a soft, learnable way. Our method only requires $\{12,24\}$ examples and has negligible computation overhead over $4$-shot ICL at inference.
@@
Components interpretation studies the function of different components in a trained model Elhage et al. (2021); Shah et al. (2024), where components could be neurons Radford et al. (2017); Wang et al. (2022a); Gurnee et al. (2023), attention heads Olsson et al. (2022), and MLPs Geva et al. (2021). To analyze the components, probing Alain and Bengio (2017), knockout Geva et al. (2023); Chang et al. (2024); Li et al. (2023), patching Wang et al. (2023a); Goldowsky-Dill et al. (2023), and early decoding nostalgebraist (2020); Geva et al. (2022) are widely used techniques. For example, Li et al. (2024) train a linear probe on every attention head to discover the truthful heads inside LLMs. Michel et al. (2019) and Voita et al. (2019) prune away a large percentage of attention heads and show that only a few are critical to the performance. Hendel et al. (2023), Liu et al. (2023), Merullo et al. (2024a), and Todd et al. (2024) view ICL as compressing demonstrations into function vectors, where they remove the demonstrations and modify (patch) the LLM activations at certain layers with the function vectors at test time. Early decoding interprets the investigated components in the textual space by projecting them through the output embedding matrix Geva et al. (2022). Our model decomposition is based on early decoding and we share some similarities with prior work Yu et al. (2023); Wang et al. (2023c), especially in discovering individual components that perform well on a task. Our contributions lie in providing a new view of ICL by decomposition, which reveals the transferability of components across diverse ICL settings.
Our Method vs. Pruning.
Our method caches the direct contributions of components to the outputs through the residual stream, i.e., logits = $\sum_{j}\mathbf{g}_{j}$. Thus, removing $\mathbf{g}_{j}$, the direct contribution of the component $j$, does not alter the contributions of the other components. In comparison, pruning a component changes the activations of the other components in later layers. In A.7, we show that pruning the good-performing components identified by our method greatly hurts the accuracy, meaning that pruning also defines these components as important Michel et al. (2019).
@@ 9 Limitations
Our component reweighting method requires a small set of labeled data $\mathcal{D}_{\text{train}}$ to train the component weights $w$. However, we believe it is not unreasonable to have at least $K=12$ labeled examples in total and we compare with baselines using the same $K$ examples. On the other hand, we do not compare with fine-tuning-based baselines, such as LM-BFF Gao et al. (2021), T-few Liu et al. (2022a), and LoRA Hu et al. (2021), because they usually require a larger GPU memory for training and more sophisticated early stopping criteria to prevent overfitting on few-shot examples. Another limitation is that we only experiment with classification tasks for ease of evaluation. We leave it for future work to generalize our method to generation tasks by doing decomposition and reweighting at every token during generation.
Despite similarities in model decomposition, the focus of this paper is not circuits in LLMs Wang et al. (2023a). Thus, we only have limited experiments towards mechanistic understanding of the curious components in section 3.3 and A.7. Unlike prior work that uses synthetic tasks to testify whether a head attends to certain tokens Dutta et al. (2024); Merullo et al. (2024b), we work on standard NLP benchmark datasets without obviously correct or incorrect tokens to collect answers, making mechanistic interpretation more challenging.
@@ Acknowledgements
We thank Johnny Wei for his valuable suggestions on the paper structure. We thank Qinyuan Ye, Ryan Wang, Gustavo Lucas Carvalho, Ameya Godbole, Wang Zhu, Daniel Firebanks-Quevedo, and the anonymous reviewers for their helpful feedback. This work was funded in part by gifts from Open Philanthropy and Cisco Research, and was also supported in part by the National Science Foundation under Grant No. IIS-2403436. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ A.1 Tasks and Templates
Table 11 summarizes the 13 datasets we use in the paper, where we construct balanced test sets by randomly sampling 2000 examples in each task. We form the prompts by concatenating demonstrations in a randomly shuffled order. To avoid the recency bias Zhao et al. (2021), we keep shuffling the demonstrations until the last two have different labels. For minimally conservative templates (section 4.1), Table 10 compares the contrast sets we construct on Llama-2-7B. For our case study on Task069 and Task070, we sample 3 templates from Sclar et al. (2024). Figure 7 compare the prompts of Task069 and Task070, which consist of an instruction followed by $K$ templated demonstrations. Originally, the two tasks have $\sim 4\%$ of parallel examples. To make our task transfer challenging, we discard these overlapped examples.